OpenStudy (anonymous):

I don't want a direct answer. I want someone to help me understand the problem, and how to get to the answer. Don't give out the answer please, thanks! Esmeralda is graphing a polynomial function as a parabola. Before she begins graphing it, explain how to find the vertex. Make sure you include how to determine if it will be a maximum or minimum point. Use an example quadratic function to help you explain and provide its graph. Thanks for helping!

OpenStudy (anonymous):

@zepdrix @mathmale @mathstudent55 @ganeshie8 @sleepyjess

OpenStudy (sleepyjess):

First, I think they mean a quadratic function...

OpenStudy (sleepyjess):

Second, right now I wish I had brought my math book with my across the country, let me see what I can find

OpenStudy (anonymous):

K

OpenStudy (sleepyjess):

Are you familiar with the form y = a(x - h)\(^2\) + k

OpenStudy (anonymous):

No I know y = mx + b :)

OpenStudy (anonymous):

Like a swag boss

OpenStudy (sleepyjess):

Hmm... that is for the equation of a line. The lesson doesn't cover anything about the equation of a parabola?

OpenStudy (anonymous):

The lesson is kinda sucky. It's not compatible with my brain. You @zepdrix @mathmale @mathstudent55 @ganeshie8 and a few others are lol :)

OpenStudy (sleepyjess):

haha :) Okay, I'll kinda have to give an answer to explain it, is that okay?

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

Okay, this is kinda off topic, but why does it say "high" on the bottom right of my question?

OpenStudy (anonymous):

@sleepyjess You've been typing for a while. Where you in the middle of typing when you had a heart attack. ARE YOU OKAY!!!!!

OpenStudy (anonymous):

@satellite73

OpenStudy (sleepyjess):

To find the vertex of a parabola, you'll need the equation in vertex form first. Vertex form is y = a(x - h)\(^2\) + k, with h being the x value of the vertex, and k being the y value of the vertex. For example, if we have y = 2(x - 4)\(^2\) + 6, 4 is the h value, and 6 is the k value, making the vertex (4, 6). Now sometimes, we'll have an equation that looks like y = \(2x^2\) + 6x + 8. This is a little trickier. There's a little trick for equations like that. We need to use h = \(\dfrac{-b}{2a}\) and k = f(h). f(h) simply means put in what we found for h into the original equation (\(2x^2\)+ 6x + 8 Let me show you: \(2x^2\) + 6x + 8 h = \(\dfrac{-6}{2(2)}\) \(\dfrac{-6}{4}\) h = -\(\dfrac{3}2\) now plug that in for x \(2(-\dfrac 32) ^2\) + 6(\(-\dfrac 32\)) + 8 -3 - 9 + 8 k = -4 so the vertex for that equation is (-\(\dfrac 32\), -4)

OpenStudy (sleepyjess):

Hopefully that all makes sense, if not I'd be happy to explain more :)

OpenStudy (anonymous):

@sleepyjess It does make sense. So basically, all you need to do is know what k and h are to find the x and y coordinates. And if nesecarry, use the formula -b/2a to find the x coordinate and f(h) to find the y, correct?

OpenStudy (sleepyjess):

Yep :)

OpenStudy (sleepyjess):

You're a quick learner :3

OpenStudy (anonymous):

@zepdrix Made me a little familiar with the topic a while ago. Thanks for helping on this @sleepyjess UR AWESOME!

OpenStudy (anonymous):

@sleepyjess This is off topic, but how many medals do you have?

OpenStudy (anonymous):

Like smartcents lol

OpenStudy (sleepyjess):

Oh yeah, the "High" means that when you signed up, the site asked you a question something like, what level will your questions be? Middle, High, College, Grad, etc. YOu must have selected high

OpenStudy (sleepyjess):

Smartcents I have 1310 :)

OpenStudy (sleepyjess):

Or if you were on the site before the update, it asked you the question when you logged on one day

OpenStudy (anonymous):

@sleepyjess Oh, I was wondering what it meant lol. 1310?!?!??!?!?!? I only have 18 :(

OpenStudy (sleepyjess):

I've been on the site for 2 years and 1 day :)

OpenStudy (anonymous):

o, that explains it. lol I want a cool title, like human calculator :( Anyways, I should probably answer this question, g2g Oh, and thanks again for helping @sleepyjess !

OpenStudy (sleepyjess):

No problem :)

OpenStudy (anonymous):

@sleepyjess Wait, how do I tell if it will be a minimum or a maximum?

OpenStudy (sleepyjess):

If "a" in the equation is positive, it opens upward, meaning the vertex will be a minimum. If "a" is negative, it opens downward, meaning the vertex will be the maximum